# Adaptive Accountability in Networked MAS:

### v1

@@ 3.3 Operational Assumptions and Constraints
Regulatory and ethical context: Many target applications fall under the EU AI Act’s “high-risk” category. Each ledger entry therefore carries a pseudonymised agent ID, a purpose-limited data tag, and a zero-knowledge proof attesting that the log originates from certified software. This design satisfies GDPR data-minimisation requirements yet still lets external auditors reconstruct causal chains. Intervention records include plain-language rationales, effect sizes, and parameter settings so that policymakers and domain experts can review system behavior without inspecting source code.
These deliberately pessimistic assumptions, i.e., lossy sensing, tight resource budgets, heterogeneous agents, and stringent legal scrutiny, shape every design choice in Sections 47. By proving guarantees and demonstrating empirical performance under such constraints, we strengthen confidence that the Adaptive Accountability Framework will generalise to real-world MAS whose operating conditions may be even harsher than those evaluated here.
4 Proposed Framework: Adaptive Accountability in Networked MAS
@@ 4.1 Architecture Overview
Figure 1 illustrates the stack, organised into four layers: agent, monitor, ledger, and governor plus a cross-cutting security plane.
@@
3. Ledger layer: Gossip streams converge at edge servers that execute Merkle-DAG synthesis: new records extend a Merkle prefix tree whose leaves reference prior sub-roots, forming a snapshotted Merkle DAG every $H_{\mathrm{snap}}=256$ steps [merkle1989certified]. The root hash of each snapshot is signed with a rotating ECDSA key and broadcast back to agents for self-verification. The DAG design, adapted from block-lattice data stores, provides $O(\log V)$ proof-of-inclusion while avoiding the confirmation latency of Nakamoto consensus.
@@ 4.3 Online Detection of Harmful Norms
Initialise : $S\leftarrow 0$; $h\leftarrow h_{0}$; gain schedule $\eta_{t}=t^{-0.6}$
@@ 13 $h\leftarrow h+\eta_{t}(\mathsf{alert}_{t}-\alpha)$ // Robbins--Monro update
In Algorithm 2, lines 1–2 accumulate evidence that $Z_{t}$ has shifted upward so whenever the statistic $S$ crosses the adaptive threshold $h$, an alarm is fired and $S$ is reset (classical Page’s rule [page1954continuous]). Lines 3–4 perform a Robbins–Monro stochastic approximation [robbins1951stochastic]: if the empirical alarm frequency exceeds the target $\alpha$ the threshold is gently raised, and vice-versa. Section 5.3 proves that this update drives the long-run false-alarm rate to $\alpha$ regardless of slow concept drift (Assumption A3).
@@
Across 360 Monte-Carlo seeds (Section 7) the combined detector recognized emergent collusion within 9\!9 steps median (95th percentile =17=\!17) and maintained a 4.7 % system-wide false-positive rate—even with 20 % record loss and bounded 3-step delays (Assumption A1). The compute footprint is negligible: updating all $|\Phi|\!=\!3$ norms costs $0.07\pm 0.01$ ms on a 2.2 GHz Xeon Silver and <0.3<\!0.3 ms on a Cortex-A55, comfortably within the five-per-cent control-loop budget set in Section 6.
@@ 4.4 Adaptive Intervention Mechanisms
The supervisory spend per step is bounded by $\mathbb{E}[\text{Cost}(\sigma_{t})]\leq\lambda+c_{\mathrm{patch}}\alpha+c_{\mathrm{throttle}}\alpha^{2}$, where $\alpha$ is the false-alarm budget and the last two terms account for patch and throttle operations that occur only on alerts. Since $\lambda>g_{\max}$ by design, Theorem 5 ensures $\limsup_{T\to\infty}C_{T}/T\leq\eta^{\star}$. In practice we observed $\eta^{\star}=0.19$ under a 5 % Byzantine mix, dropping to 0.06 in fully rational populations (§7).
@@ 4.5 Scalability and Fault Tolerance
At runtime, the ledger is materialized as a Merkle directed-acyclic graph (Merkle DAG) whose nodes are addressed by the cryptographic hash of their content [merkle1989certified]. Every 256-step snapshot is edge-partitionable: we shard by the first 16 bits of the event hash $h_{e}$, creating $2^{16}$ disjoint key ranges that can be balanced across the edge tier. Each shard is replicated onto f+1f{+}1 servers; read and write clients contact a quorum, i.e., a strict majority of replicas, so the system tolerates $f$ crash failures without sacrificing linearizable semantics [gifford1979weighted]. Any two quorums overlap in at least one live replica, eliminating the need for a global lock or leader election.
@@ 4.6 Lifecycle Integration
In online adaptation, after go-live, CUSUM thresholds $h_{t}$ are tuned by mirror descent, ECDSA [^1] Elliptic Curve Digital Signature Algorithm [johnson2001ecdsastandard]. keys rotate every 24h, and new norms can be hot-enabled without downtime because the ledger schema is append-only and therefore backward-compatible.
@@ 5 Theoretical Analysis and Guarantees
This section formalises the statistical and game–theoretic properties that underpin the Adaptive Accountability Framework (AAF). We proceed in four logical stages:
@@ 5.1 Preliminaries and Standing Assumptions
We work on a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ where $\mathcal{F}_{t}$ is generated by the global MAS trajectory up to time $t$ [hernandez2019survey]. Indicators, expectations, and probabilities are taken with respect to $\mathbb{P}$ unless stated otherwise. Throughout we invoke four assumptions—justified empirically in Section 3.3—to isolate the essential trade-offs:
@@
Budgeted adversary: Byzantine agents harvest at most $g_{\max}$ expected utility per step; the supervisor’s cost budget honours $c_{\max}>g_{\max}$.
@@ 5.2 Ledger Soundness
A reader may wonder: What exactly is stored in the audit ledger, how are causal claims derived, and why do responsibility scores behave like probabilities? This subsection therefore proceeds in three steps. We (i) introduce precise definitions for the ledger structures and the responsibility rule, (ii) prove that those scores are well-normalised and converge even when packets are lost, and (iii) establish a time-uniform bound on spurious causal edges. The notation follows Section 3.1; assumptions A1–A3 remain in force.
@@
The ledger DAG gives a tamper-evident who-influenced-what map; the $\beta^{|p|}$ discount keeps distant, low-credence influences from dominating local, high-credence ones. The next lemma shows the rule is well-normalised, so $\rho_{i}(e)$ can be treated probabilistically.
Lemma 1 (Normalisation)
@@
Apply the Chernoff–Ville inequality to the self-normalised stream of F-statistics and union-bound over time [howard2021time]. Full derivation in Appendix A.3.
@@ 5.3 Detector Reliability
    the detection delay grows at most linearly with the decision threshold, matching the classical Page–Lorden benchmark [page1954continuous, lorden1971procedures].
@@ A. False-alarm control:
where $h^{\star}=\lim_{t\to\infty}h_{t}$. The bound is identical to the classical Lorden delay [lorden1971procedures] except that $h^{\star}$ is learned online rather than fixed a priori.
@@ 5.4 Intervention Efficacy
  • Penalty amplitude $\lambda$: Raising $\lambda$ tightens the compromise ceiling $\eta^{\star}$ but increases per-step supervisory cost and the risk of over-penalising honest but noisy agents.
@@
Re-arrange $\eta^{\star}=\alpha H/(\lambda H-g_{\max})$ for $\lambda$.
@@
The intervention rule can be viewed as a potential-based reward shaper [ng1999policy] where the shaping term is the scaled responsibility score. Potential-based shaping is known to preserve the set of optimal policies in single-agent MDPs; our theorem extends the intuition to multi-agent settings with partial observability by showing that undesirable equilibria become transient once shaping makes them strictly sub-optimal.
@@
In Section 7 we instantiate $\lambda=\lambda_{\min}+0.1$ and observe $\eta_{\text{emp}}=0.06\pm 0.01$ in a population of rational learners and $\eta_{\text{emp}}=0.19\pm 0.03$ when 5% of agents are Byzantine, matching the theoretical ceiling to within one standard error.
@@ 5.5 Resource Complexity
  • Storage: $T(N+d_{\max}h)\times 40\,\text{B}\approx 492\,\text{MB}$ for the entire Merkle DAG—only 3% of the 16 GB RAM in our commodity edge boxes.
  • Bandwidth: $(N+d_{\max}h)\times 40\,\text{B}\approx 80\,\text{KB s}^{-1}$, negligible relative to a 1 Gbit s-1 NIC.
  • CPU: Merkle insertion is $O(\log|\mathcal{V}_{T}|)$ with a small constant (prefix compression); incremental Granger updates are $O(m)$ with $m=8$. Profiling on an EPYC 7742 shows a mean of $87\pm 9\,\mu\text{s}$ CPU per event, i.e. $1.5\%$ of a single core at 50 Hz control frequency.
@@
Together with Theorems 35, the above bounds confirm that AAF achieves auditable accountability at bounded cost: statistical guarantees hold under packet loss, policy drift, and moderate Byzantine behavior, while the resource footprint fits comfortably within off-the-shelf edge hardware. Section 7 corroborates the analytic predictions across 360 Monte-Carlo trials and performs sensitivity tests over $\varepsilon_{\text{loss}}$, $g_{\max}$, and graph topology, demonstrating that the cost envelope remains stable under realistic operational variability.
@@ 6 Implementation Details
Section 4 described the design logic of the Adaptive Accountability Framework (AAF). Here we present the systems engineering choices that turn that logic into executable, reproducible artefacts. We walk from the lowest-level event record up to dashboards and DevOps pipelines, so readers can (i) re-run every experiment in Section 7, (ii) port AAF to their own MARL stack, and (iii) audit performance or security claims without inspecting privileged code. Throughout, we respect three engineering constraints: per-agent overhead <<5% of the control loop, strict library agnosticism, and regulator-grade provenance.
6.1 Technical Stack and Runtime Environment
Languages and libraries:
@@ 6.1 Technical Stack and Runtime Environment
Hardware profile:
All experiments (§7) run on a Slurm cluster: dual-socket AMD EPYC-7742 nodes (64 c $\times$ 2, 256 GiB RAM) plus one NVIDIA A40 GPU per job when deep RL is enabled. The CPU governor is fixed to “performance”; Linux 6.7-LTS runs with the complete-preemption patch to minimize scheduler noise.
Containerization:
Every component ships as an OCI image built via nix-flakes. Exact wheel versions and OS packages are pinned, and the resulting SHA-256 digest is embedded into each ledger snapshot—useful when an auditor needs to prove reproducibility long after deployment.
6.2 Ledger Construction and Storage Back-End
Each control step yields a 40-byte record
$e=\langle t,\,i,\,\textsf{BLAKE3}(o_{i,t})_{16},\,\textsf{BLAKE3}(a_{i,t})_{16},\,r_{i,t}\rangle,$
[…]
6.3 Message Bus, Clocks, and Failure Semantics
Dual-plane networking.
The control plane (heartbeats, digests, alarms) rides a tri-replicated Apache Kafka 3.7 cluster with exactly-once semantics. The data plane (high-rate state broadcasts) uses ZeroMQ PUB/SUB encrypted by CurveCP; a 128-entry loss-tolerant queue cushions transient link outages. If a subscriber lags >>128 messages it issues I_NEED_CATCHUP and replays the gap from a Kafka compacted topic.
Logical time.
Each record carries a Hybrid Logical Clock (HLC) $\langle\text{wall},\text{ctr}\rangle$ that merges a 48-bit physical nanosecond time-stamp with a 16-bit Lamport counter, preserving causal order even under ±250 ms NTP skew [kulkarni2014logical]. In a 100-node WAN emulation the worst-case mis-ordering was 0.8 ms.
Fault model.
All gRPC calls are retried with exponential back-off. A dual circuit breaker in envoy-proxy halts cascading failure: agent-side breakers trip at >>20% error rate; server-side breakers trip at P50 latency >>75 ms. Agents then fall back to local logging and best-effort UDP gossip until the breaker half-opens.
6.4 Detector Runtime and Intervention Delivery
Supervisor replicas subscribe to FoundationDB deltas every 20 ms via a watch. An asyncio task (i) updates norm-specific $Z^{(m)}_{t}$ vectors, (ii) evolves the CUSUM state $S^{(m)}_{t}$, and (iii) writes $\langle S_{t},h_{t}\rangle$ to Redis-5 with TTL = 30 s. Mean compute cost is 0.07 ms per cycle on a 2.2 GHz Xeon—<<1% of the control loop.
When an alarm fires the governor publishes a RewardPatch Protobuf over ZeroMQ. Agents verify an ECDSA-P-384 signature, apply the penalty in $\mathcal{O}(1)$, and—if running a neural policy—hot-load a two-layer “safety head” compiled to WebAssembly using wasmtime-17. P99 alarm-to-patch latency is <<22 ms on a 1 Gbit LAN.
6.5 Security and Compliance Tooling
Cryptography.
Event digests use keyed BLAKE3; hybrid TLS 1.3 sessions negotiate x25519-Kyber768 for post-quantum forward secrecy [bos2018crystals]. Snapshot hashes are counter-signed in an HSM with ECDSA-P-384 and time-stamped via RFC-3161.
PII minimisation.
Agent IDs are pseudonymised via AES-SIV; linking records requires access to both the ledger and an escrowed look-up table, meeting the “unlinkability unless audited” clause of the EU AI Act.
6.6 DevOps, CI/CD, and Observability
Nightly Monte-Carlo sweeps (10 seeds, 5 topologies, 3 adversary mixes) run in GitHub Actions via act-runners. Latency or FP-rate regressions block merge to main. Production deploys via Helm on Kubernetes 1.29 (Calico CNI); canary replicas (10% traffic shadow) must sustain FP < 5% and patch delivery < 50 ms for 15 min or Argo rolls back automatically. Prometheus 2.51 scrapes 87 metrics per pod; Grafana panels visualize (i) long-run compromise ratio, (ii) per-agent SHAP responsibility, and (iii) alarm-to-patch latency histogram. PDF snapshots ship weekly to a WORM-locked S3 bucket.
6.7 User-Facing Interfaces
Operational dashboard.
A Next.js 14 front-end (OIDC via Keycloak) renders social reward, Gini index, active alerts, and a DAG explorer that reconstructs the causal sub-graph around sentinel event $e^{\star}$; edges are color-coded by responsibility mass.
Data-science notebook.
A multi-tenant JupyterHub mounts read-only FoundationDB snapshots via fdb-sqlean. Analysts can run SQL-like queries, export to Pandas, and push results into statistical notebooks. Role-based access control enforces “least privilege” (GDPR Recital 78).
6.8 Portability and Edge Deployment
AAF touches only the message bus (digest gossip) and the reward API. Any MARL stack that exposes these hooks—RLlib, PettingZoo, OpenSpiel, Unity ML-Agents—can host AAF. For resource-starved hardware we supply a 32-KB Rust micro-kernel that maintains the ring buffer, computes BLAKE3 hashes, and emits CAN-FD frames on an ARM Cortex-M7 running at 600 kHz.
With these engineering choices AAF remains bandwidth-frugal, cryptographically auditable, and platform-agnostic, fully instantiating the theoretical model of Sections 45. The next section benchmarks this implementation and probes how closely empirical behavior tracks our analytical bounds.
7 Experimental Setup and Case Studies
This section empirically evaluates the Adaptive Accountability Framework (AAF). We (i) describe the simulation environment and parameter grid, (ii) state the evaluation metrics and baselines, and (iii) report quantitative and qualitative results that validate the analytical guarantees of Section 5. All experiments are fully reproducible: a one-line nix flake command (‘nix run github:alqithami/aaf-sim’) launches the end-to-end pipeline on either local CPU or the Slurm cluster described in §6.
7.1 Simulation Environment
Our principal benchmark is a resource-sharing game that abstracts distributed power-grid dispatch, shared wireless bandwidth, or cloud capacity brokering. The environment maintains a pool of $R_{\max}=100$ resource units replenished at a fixed rate $R_{\text{in}}$ per step. An agent $A_{i}$ submits a request $q_{i,t}\in[0,q_{\max}]$ and receives an allocation $a_{i,t}$ from a proportional–share rule modulated by the distribution policy parameter:
$a_{i,t}=\begin{cases}\displaystyle q_{i,t}&\text{if }\sum_{j}q_{j,t}\leq R_{\max},\\[6.0pt] \displaystyle\frac{q_{i,t}^{\alpha}}{\sum_{j}q_{j,t}^{\alpha}}\,R_{\max}&\text{otherwise},\end{cases}$
where $\alpha=1$ corresponds to fair share and $\alpha\to 0$ enforces strict proportionality; dist_policy=none sets $\alpha\!=\!1$. Agents observe only their own allocation plus a noisy 2-hop neighborhood view on a Watts–Strogatz graph [watts1998collective]; partial observability is toggled by exposing (on) or hiding (off) a noisy global queue length.
Each agent earns $r_{i,t}^{\text{private}}=a_{i,t}-\text{penaltyFactor}\cdot\mathbf{1}\{q_{i,t}\geq 0.6R_{\max}\}$, so large greedy requests incur a cost. A SOCIAL bonus $r_{t}^{\text{social}}=\frac{1}{N}\sum_{i}a_{i,t}$ is combined with weight $\lambda=0.3$. The penalty term is exactly what the governor modulates during interventions (§4.4).
7.2 Experimental Design and Parameters
Monte-Carlo grid.
We sweep the Cartesian product of
  • Population size: $N\in\{10,50,100\}$.
  • Penalty factor: $\{0.05,\,0.20,\,0.35\}$.
  • Partial observability: on vs. off.
  • Distribution policy: $\alpha\in\{0\,(\texttt{none}),\,0.25,\,1.0\}$.
With five random seeds per cell the design comprises $3\times 3\times 2\times 3\times 5=270$ full-length runs (1,000 steps each). An additional $90$ runs inject a 5% Byzantine sub-population that always requests the maximum resource.
Baselines.
  1. 1.
    PPO‐only: agents train under proximal policy optimization without any audit or intervention; this is the strongest “naïve learning” baseline.
  2. 2.
    Static guard: a centralized threshold caps $q_{i,t}\leq 0.6R_{\max}$ but performs no causal tracing.
Evaluation metrics.
(i) Norm-violation ratio $C_{T}/T$; (ii) Detection delay $T_{\text{alarm}}-\tau^{\star}$; (iii) Social welfare $J_{\text{soc}}=\tfrac{1}{T}\sum_{t}\tfrac{1}{N}\sum_{i}r_{i,t}$; (iv) Inequality via the Gini index [gini1912variabilita]; and (v) Bandwidth overhead.

### v2

@@ 3.3 Operational Assumptions and Constraints
Regulatory and ethical context: Many target applications fall under the EU AI Act’s “high-risk” category. Each ledger entry therefore carries a pseudonymized agent ID, a purpose-limited data tag, and a zero-knowledge proof attesting that the log originates from certified software. This design satisfies GDPR data-minimization requirements yet still lets external auditors reconstruct causal chains. Intervention records include plain-language rationales, effect sizes, and parameter settings so that policymakers and domain experts can review system behavior without inspecting source code.
These deliberately pessimistic assumptions, i.e., lossy sensing, tight resource budgets, heterogeneous agents, and stringent legal scrutiny, shape every design choice in Sections 47. By proving guarantees and demonstrating empirical performance under such constraints, we strengthen confidence that the Adaptive Accountability Framework will generalize to real-world MAS whose operating conditions may be even harsher than those evaluated here.
4 Proposed Framework: Adaptive Accountability in Networked Multi-Agent Systems
@@ 4.1 Architecture Overview
Figure 1 illustrates the stack, organized into four layers: agent, monitor, ledger, and governor plus a cross-cutting security plane.
@@
3. Ledger layer: Gossip streams converge at edge servers that execute Merkle-DAG synthesis: new records extend a Merkle prefix tree whose leaves reference prior sub-roots, forming a snapshotted Merkle DAG every $H_{\mathrm{snap}}=256$ steps [36]. The root hash of each snapshot is signed with a rotating ECDSA key and broadcast back to agents for self-verification. The DAG design, adapted from block-lattice data stores, provides $O(\log V)$ proof-of-inclusion while avoiding the confirmation latency of Nakamoto consensus.
@@ 4.3 Online Detection of Harmful Norms
Initialize : $S\leftarrow 0$; $h\leftarrow h_{0}$; gain schedule $\eta_{t}=t^{-0.6}$
@@ 13 $h\leftarrow h+\eta_{t}(\mathsf{alert}_{t}-\alpha)$ // Robbins--Monro update
In Algorithm 2, lines 1–2 accumulate evidence that $Z_{t}$ has shifted upward so whenever the statistic $S$ crosses the adaptive threshold $h$, an alarm is fired and $S$ is reset (classical Page’s rule [46]). Lines 3–4 perform a Robbins–Monro stochastic approximation [55]: if the empirical alarm frequency exceeds the target $\alpha$ the threshold is gently raised, and vice-versa. Section 5.3 proves that this update drives the long-run false-alarm rate to $\alpha$ regardless of slow concept drift (Assumption A3).
@@ 4.4 Adaptive Intervention Mechanisms
The supervisory spend per step is bounded by $\mathbb{E}[\text{Cost}(\sigma_{t})]\leq\lambda+c_{\mathrm{patch}}\alpha+c_{\mathrm{throttle}}\alpha^{2}$, where $\alpha$ is the false-alarm budget and the last two terms account for patch and throttle operations that occur only on alerts. Since $\lambda>g_{\max}$ by design, Theorem 5 ensures $\limsup_{T\to\infty}C_{T}/T\leq\eta^{\star}$. In §7 we evaluate this ceiling empirically; across a broad grid of regimes, AAF reduces executed compromise relative to PPO while preserving welfare.
@@ 4.5 Scalability and Fault Tolerance
At runtime, the ledger is materialized as a Merkle directed-acyclic graph (Merkle DAG) whose nodes are addressed by the cryptographic hash of their content [36]. Every 256-step snapshot is edge-partitionable: we shard by the first 16 bits of the event hash $h_{e}$, creating $2^{16}$ disjoint key ranges that can be balanced across the edge tier. Each shard is replicated onto f+1f{+}1 servers; read and write clients contact a quorum, i.e., a strict majority of replicas, so the system tolerates $f$ crash failures without sacrificing linearizable semantics [21]. Any two quorums overlap in at least one live replica, eliminating the need for a global lock or leader election.
@@ 4.6 Lifecycle Integration
In online adaptation, after go-live, CUSUM thresholds $h_{t}$ are tuned by mirror descent, ECDSA [^1] Elliptic Curve Digital Signature Algorithm [29]. keys rotate every 24h, and new norms can be hot-enabled without downtime because the ledger schema is append-only and therefore backward-compatible.
@@ 5 Theoretical Analysis and Guarantees
This section formalizes the statistical and game–theoretic properties that underpin the Adaptive Accountability Framework (AAF). We proceed in four logical stages:
@@ 5.1 Preliminaries and Standing Assumptions
We work on a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ where $\mathcal{F}_{t}$ is generated by the global MAS trajectory up to time $t$ [24]. Indicators, expectations, and probabilities are taken with respect to $\mathbb{P}$ unless stated otherwise. Throughout we invoke four assumptions—justified empirically in Section 3.3—to isolate the essential trade-offs:
@@
Budgeted adversary: Byzantine agents harvest at most $g_{\max}$ expected utility per step; the supervisor’s cost budget honors $c_{\max}>g_{\max}$.
@@ 5.2 Ledger Soundness
A reader may wonder: What exactly is stored in the audit ledger, how are causal claims derived, and why do responsibility scores behave like probabilities? This subsection therefore proceeds in three steps. We (i) introduce precise definitions for the ledger structures and the responsibility rule, (ii) prove that those scores are well-normalized and converge even when packets are lost, and (iii) establish a time-uniform bound on spurious causal edges. The notation follows Section 3.1; assumptions A1–A3 remain in force.
@@
The ledger DAG gives a tamper-evident who-influenced-what map; the $\beta^{|p|}$ discount keeps distant, low-credence influences from dominating local, high-credence ones. The next lemma shows the rule is well-normalized, so $\rho_{i}(e)$ can be treated probabilistically.
Lemma 1 (Normalization)
@@
Apply the Chernoff–Ville inequality to the self-normalized stream of F-statistics and union-bound over time [27]. Full derivation in Appendix A.3.
@@ 5.3 Detector Reliability
    the detection delay grows at most linearly with the decision threshold, matching the classical Page–Lorden benchmark [46, 34].
@@ A. False-alarm control:
where $h^{\star}=\lim_{t\to\infty}h_{t}$. The bound is identical to the classical Lorden delay [34] except that $h^{\star}$ is learned online rather than fixed a priori.
@@ 5.4 Intervention Efficacy
  • Penalty amplitude $\lambda$: Raising $\lambda$ tightens the compromise ceiling $\eta^{\star}$ but increases per-step supervisory cost and the risk of over-penalizing honest but noisy agents.
@@
Rearrange $\eta^{\star}=\alpha H/(\lambda H-g_{\max})$ for $\lambda$.
@@
The intervention rule can be viewed as a potential-based reward shaper [43] where the shaping term is the scaled responsibility score. Potential-based shaping is known to preserve the set of optimal policies in single-agent MDPs; our theorem extends the intuition to multi-agent settings with partial observability by showing that undesirable equilibria become transient once shaping makes them strictly sub-optimal.
@@
In Section  efsec:experiments we empirically validate that the observed compromise ratios remain bounded and that AAF consistently improves the compromise–welfare trade-off across the full parameter sweep.
@@ 5.5 Resource Complexity
  • Storage (order-of-magnitude): assuming 40 B per event record, the log volume is on the order of $T(N+d_{\max}h)\times 40$ B; for $\langle N\!=\!100,\;h\!=\!8,\;d_{\max}\!\leq\!8,\;T\!=\!10^{6}\rangle$ this is $\approx 492\,\text{MB}$.
  • Bandwidth (order-of-magnitude): per-step gossip is $O((N+d_{\max}h)\times 40)$ B; at 50 Hz this corresponds to $\approx 80\,\text{KB\,s}^{-1}$ under the same parameterization.
  • Compute: the per-step detector and attribution updates are linear in the local window $h$ with small constants; we record wall-clock runtimes in the exported summaries for completeness.
@@
Together with Theorems 35, the above bounds confirm that AAF achieves auditable accountability at bounded cost: statistical guarantees hold under policy drift and moderate Byzantine behavior, while the resource footprint is linear in $N$ and $T$. Section 7 corroborates these analytic predictions across 87,48087{,}480 simulation runs, sweeping population size, penalty_factor, redistribution $\alpha$, observability, and Byzantine rates.
@@ 6 Implementation Details
We provide a self-contained simulation artifact that implements the Adaptive Accountability Framework (AAF) end-to-end and reproduces all tables and figures in §7. The implementation is intentionally lightweight: all components are written in Python and run on either CPU or a single GPU (optional) using PyTorch for the learning baselines.
6.1 Code organization
The artifact is organized as a small Python package (aaf_q1/) plus a set of reproducibility scripts (scripts/):
  • aaf_q1/envs: the resource_sharing and public_goods environments.
  • aaf_q1/agents: PPO and baseline variants.
  • aaf_q1/aaf: detector, attribution, and intervention modules.
  • scripts/make_grid.py: generate JSONL experiment grids.
  • scripts/run_grid.py: execute grids with multiprocessing and optional sharding.
  • scripts/aggregate.py: materialize analysis/all_runs_flat.csv and analysis/final_summary.csv.
  • scripts/make_latex.py and scripts/make_figures.py: produce paper-ready tables and figures.
  • scripts/stats.py: seed-matched paired tests.
6.2 Logging
Each run writes a config.json (exact hyperparameters) and a summary.json (scalar metrics). When log_mode=steps, we additionally export step_logs.csv (time-series traces used for learning-curve plots).
6.3 Reproducibility
The full pipeline is driven by explicit command-line scripts; the repository README contains the exact commands used to produce the results in this manuscript, including sharded execution for multi-node runs. Our aggregated results and statistical tests are derived only from the exported CSV/JSON logs, so third parties can re-run the analysis without access to hidden state.
6.4 Mapping to deployment
The full AAF design in §4 assumes an append-only audit ledger and trustworthy message delivery. In our simulation artifact we model this ledger as an append-only event log and evaluate the statistical (detection/attribution) and control (intervention) components. Cryptographic signing and distributed storage are orthogonal engineering choices and can be layered into a deployment without changing the learning or detection logic.
7 Experimental Setup and Case Studies
This section empirically evaluates the Adaptive Accountability Framework (AAF). We (i) describe the simulation environment and parameter grid, (ii) state the evaluation metrics and baselines, and (iii) report quantitative and qualitative results that validate the analytical guarantees of Section 5. All experiments are fully reproducible using the released artifact: we provide explicit scripts to generate the experiment grid, execute runs, aggregate metrics, and regenerate all tables and figures (see the repository README at https://github.com/alqithami/AAF).
7.1 Simulation Environment
We evaluate AAF in two stylized multi-agent benchmarks that exhibit emergent norms under self-interested learning and admit clear notions of “policy compromise”.
Resource-sharing game.
Agents repeatedly request a divisible resource subject to a shared capacity $C$. At time $t$, each agent $i$ proposes a requested quantity $q_{i,t}\in[0,r_{\max}]$; an environment scheduler executes $q^{\mathrm{exec}}_{i,t}$ such that $\sum_{i}q^{\mathrm{exec}}_{i,t}\leq C$. We flag a norm violation (“greedy” request) when $q^{\mathrm{exec}}_{i,t}\geq\gamma r_{\max}$, and apply a penalty scaled by penalty_factor. Each agent’s reward combines private utility and a social term,
$r_{i,t}=r_{i,t}^{\text{private}}+\lambda r_{t}^{\text{social}},\qquad r_{i,t}^{\text{private}}=q^{\mathrm{exec}}_{i,t}-\texttt{penalty\_factor}\cdot\mathbb{1}[\text{violation}_{i,t}],$
where the scalar $\lambda\in[0,1]$ controls the trade-off between egoistic and collective incentives. Observations include the agent’s current allocation and local neighbourhood statistics on the communication graph; when partial_obs=1 we additionally provide a noisy scalar summary of the global queue length.
Public-goods game.
Each agent contributes $c_{i,t}\in[0,E]$ from an endowment $E$ to a common pot. All agents receive a shared return $\mathrm{ret}_{t}=m\cdot\sum_{i}c_{i,t}/N$ (multiplier $m>1$), and the private payoff is
$r_{i,t}^{\text{private}}=(E-c_{i,t})+\mathrm{ret}_{t}-\texttt{penalty\_factor}\cdot\mathbb{1}[c_{i,t}<c_{\min}],$
penalising “free-riding” below a minimum contribution $c_{\min}$. Rewards again add a social term $\lambda\cdot\mathrm{ret}_{t}$. When partial_obs=1, agents receive an additional noisy public signal of mean contribution.
Communication overhead.
Each step includes a small control-plane message (e.g., a signed action log or lightweight telemetry); AAF augments this with minimal provenance and detector metadata needed for accountability. We measure total bytes as bandwidth_overhead_bytes in our evaluation logs.
7.2 Experimental Design and Parameters
Factorial grid and seeds.
@@ 7.2 Experimental Design and Parameters
Baselines and ablations.
We compare AAF against three learning baselines and one rule-based oracle, and include four ablations to isolate each component:
  • PPO-only: standard Proximal Policy Optimization (PPO) with no accountability layer.
  • Static-guard (oracle): a non-learning safety clamp that hard-limits unsafe actions, yielding zero executed compromises by construction.
  • Constrained PPO: PPO with a Lagrangian penalty on norm-violation events (no provenance or attribution).
  • Fair PPO: PPO with a Gini regularizer on allocations.
  • AAF-full: online change-point detection (adaptive cumulative sum (CUSUM)), responsibility scoring, and interventions (reward shaping + targeted policy patching).
  • Ablations: detector-only, shaping-only, patch-only, and no-attribution.
Metrics and statistical testing.
We report (i) compromise ratio (attempted and executed), (ii) social welfare $\sum_{i}R_{i}$, (iii) allocation inequality via the Gini index, (iv) alarm count and detection delay (for methods with detectors), and (v) attribution accuracy (top-1 and recall@$k$) under Byzantine injections. For grid-level comparisons we run seed-matched paired tests (10 pairs per regime) between AAF-full and PPO-only. We report the fraction of regimes significant at Holm–Bonferroni corrected $p<0.05$ across the 324 regimes, as well as median relative effects.